# Supernovae in 2023 (review): breakthroughs by late observations

### v3

@@ 3.2 The breakthrough: evidence for exploding jets
From Section 3.1 it is evident that although the delayed neutrino explosion mechanism is more popular in the literature than the JJEM, it nonetheless encounters some challenges. Here I will argue that the point-symmetric morphologies of three CCSNRs that were studied in 2023 strongly support the JJEM and pose severe problems to the neutrino-driven explosion mechanism.
@@
Late observations, from months post-explosion before the nebular phase (e.g., Haynie & Piro 2023; Medler et al. 2023), through the nebular phase (e.g., DerKacy et al. 2023b; Dessart et al. 2023a, b; Ertini et al. 2023; Fang & Maeda 2023; Liljegren et al. 2023; Margutti et al. 2023; Omand & Jerkstrand 2023; Ravi et al. 2023; Rizzo Smith, Kochanek, & Neustadt 2023; Wesson et al. 2023; Xiang et al. 2023b, and then the remnant phase (e.g., Sato et al. 2023; Zsíros et al. 2023 and more below) can teach us about the progenitor properties and its environment at the explosion time and on the explosion mechanism. Before I describe evidence for jet-driven CCSNe by analyzing the CCSNR phase, I refer to the prediction of the delayed neutrino explosion mechanism that some massive stars collapse into a black hole with no explosion (or with only a very faint explosion). This collapse is termed ‘failed supernovae’.
@@
Some papers in 2023 continue to study this concept of a ‘failed supernova’ (e.g., Coughlin 2023; da Silva Schneider & O’Connor 2023; Disberg & Nelemans 2023; Kuroda & Shibata 2023; Li et al. 2023a; Paradiso et al. 2023). A ‘failed supernova’ is a prediction of the delayed neutrino explosion mechanism where neutrino heating cannot explode the star which in turn collapses to a black hole. On the other hand, in the JJEM there are no failed CCSNe. Even if a black hole is formed there is an explosion, and more likely an energetic explosion (see review Soker 2022). In the JJEM a black hole is likely to be formed when the pre-collapse core is rapidly rotating because this leads to fixed-axis jets that are less efficient in exploding the core (Soker 2023b), i.e., an inefficient jet feedback mechanism (for a recent paper on the accretion disk around the central black hole in CCSNe see Fujibayashi et al. 2023b). Soker (2023b) studies the continuum of explosion outcomes from zero pre-collapse core rotation to rapid pre-collapse rotation in the frame of the JJEM.
@@
The new study by Strotjohann, Ofek, & Gal-Yam (2023) and early results of Byrne & Fraser (2022) question the existence of a large population of ‘failed supernovae’. The detection of a black hole at the center of an SNR, as tentatively claimed by Balakrishnan et al. (2023) for SNR G351.9-0.9, shows that black holes can form in CCSNe, rather than in ‘failed supernovae’. The detection by Beasor et al. (2023) of a luminous infrared source at the position of the ‘failed supernova’ candidate N6946-BH1 shows, as they claim, that it better fits a type II intermediate luminosity optical transient (type II ILOT) than a ‘failed supernova’ (but see Kochanek 2023; Kochanek, Neustadt, & Stanek 2023). The above-cited studies challenge the prediction of the delayed neutrino explosion mechanism of ‘failed supernovae’. I reiterate my claim that studies of CCSNe that form black holes cannot ignore jets.
@@
Let me then turn to a prediction of the JJEM that in some (but not all) cases the jittering jets imprint a point-symmetric morphology. Most of the jittering jets during the explosion process, typically lasting for about one to several seconds and more, deposit their entire energy in the core of the star and explode it. From that point on the explosion is similar (but not identical) to that of the delayed-neutrino explosion mechanism. However, in some cases, the few last pairs of jets that are launched after the inner parts of the stars are exploded might leave an imprint on the outer ejecta. In most cases only the last pair leaves an imprint, forming two opposite ‘ears’ (e.g., Soker 2023d). An ear is a protrusion from the main ejecta with a decreasing cross-section with distance from the center. In a minority of cases, more than one pair of jets might leave imprints, forming a point-symmetric morphology.
@@
In 2023 I identified point-symmetric morphologies in three CCSNRs. In Section 1 I discussed the point-symmetric morphology and why jets are most likely to shape such morphologies.
@@
In Soker (2024d) I studied the morphology of the CCSNR N63A based on new X-ray images that Karagöz et al. (2023) present. I present the image and my identification of a point-symmetric morphology in Figure 1. I identified the point-symmetric morphology to be the three pairs of ears that are marked in the figure. This point symmetry cannot be shaped by ejecta-CSM interaction or ejecta-ISM interaction (Section 1). Therefore, this point-symmetry strongly supports the JJEM. Moreover, the differences in sizes of the two ears in each pair are accounted for by the JJEM.
@@
In each pair, one ear is significantly larger than the opposite ear. This is not due to an asymmetrical medium on sides A and B (marked in the figure) as Ear 1a is larger than Ear 1b, but Ears 2a and 3a are smaller than Ears 2b and 3b, respectively. I, therefore, concluded (Soker, 2024d) that the asymmetry of each pair results from asymmetrical opposite jets at the time of the jets’ launching. This asymmetry likely results from that the accretion disk that launches the jets in the JJEM has no time to fully relax. In the JJEM there are $\approx{\rm few}$ to $\approx 30$ jet-launching episodes, each lasting for $\approx 0.01-0.1{~{}\rm s}$. The viscous timescale of the accretion disk in each jet-launching episode is not much shorter than this timescale. Therefore, the disk has no time to fully relax to a thin accretion disk. Since the instabilities that must operate in the JJEM are likely to form each accretion disk with unequal sides of the temporarily equatorial plane, this asymmetry lasts for most of the jet-launching episode. The two unequal sides of the accretion disk are very likely to launch two opposite jets that have different energy, mass, and velocity. These asymmetrical two opposite jets in a pair will inflate unequal ears.
@@
In Soker (2023d) I classified 14 CCSNRs into five groups according to their morphology as shaped by jets. I argued that the degree of pre-collapse core rotation is one of the main parameters that determines the morphological class. The faster the pre-collapse core rotation is the less random the directions of the jets in the different jet-launching episodes are. For a very slow (or zero) pre-collapse core rotation, the jets jitter almost randomly. This, I suggested, might lead to a point-symmetric morphology with pairs of opposite structures in many directions. In that study, I identified a point-symmetric morphology in the Vela SNR. The Vela SNR is well-studied (e.g., Mayer et al. 2023; Ritchey 2023; Kargaltsev, Hare, & Lange 2024). I present the identification of the point-symmetric morphology in Figure 2 taken from Soker (2023d). Earlier studies already noted the existence of pairs of opposite structural features, as marked in the figure. However, only in Soker (2023d) did I explore the full point symmetry and connect it to jittering jets that exploded the progenitor of the Vela SNR.
@@
Figure 2: The point-symmetric morphology of the Vela SNR, as identified by Soker (2023d). The ROSAT X-ray image is from Aschenbach et al. (1995), and the figure is adapted from figure 1 of Sapienza et al. (2021). Clumps A to F were identified by Aschenbach et al. 1995. The white AG-line and the labeling of the clumps are by Sapienza et al. (2021). The thick-yellow FJ-line and DE-line and the two dashed-black lines are marked by Soker (2023d). Each of the two dashed-black lines connects a clump to an assumed counter jet.
@@
SN 1987A has been one of the prime focuses of the SN community since its explosion (for 2023 papers see, e.g., Arendt et al. 2023; Cikota et al. 2023; Dohi et al. 2023; Fiorillo et al. 2023; Jones et al. 2023; Kangas et al. 2023; Larsson et al. 2023; Weishi Li et al. 2023; Bouchet et al. 2024; Chen, Chen, & Ono 2024; Fransson et al. 2024; Ravi et al. 2024; Sapienza et al. 2023). I am interested in the ejecta morphology and its implications on the explosion mechanism. Ono et al. (2023) consider a bipolar explosion morphology for SN 1987A. They do not refer to jets, though.
@@
Figure 3: An image of SN 1987A from JWST [source: NASA Webb Telescope Team, AUG 31, 2023; NASA, ESA, CSA: Matsuura, M. (Cardiff University), Arendt, R. (NASA’s Goddard Spaceflight Center & University of Maryland, Baltimore County), Fransson, C.]. [^3] For the original image see: https://www.nasa.gov/missions/webb/webb-reveals-new-structures-within-iconic-supernova/ Red represents emission at 4.44 microns, orange at 4.05 microns, yellow at 3.23 microns, cyan at 1.64 and 2.0 microns, and blue at 1.5 microns. The arrows in the left panel are additions from Soker (2024b). These arrows point at clumps, namely, some small areas in the outer regions of the blue-cyan structure that are brighter than their surroundings. The three black lines in the right panel, also from Soker (2024b), connect pairs of clumps that the arrows in the left panel point at. The red double-headed arrow indicates the direction of the elongated axis of the blue-cyan structure. The two red-dashed lines signify a possible structure of two opposite jets.
Let me end with the study of the northeast jet of the SNR Cassiopeia A by Koo et al. (2023b). They conclude that both the neutrino-driven explosion mechanism and the magnetocentrifugal jet-induced explosion (fixed-axis jets) have problems accounting for the properties of the northeast jet of Cassiopeia A. This is a known result. As I commented in the past, the JJEM can best explain the properties of Cassiopeia A (e.g., Soker 2017). New results (after this paper’s first and second versions were posted on the arXiv) by Bear & Soker (2024) show that Cassiopeia A also has a very robust point symmetry, as expected by the JJEM. An even later paper (Soker, 2024e) identifies point symmetry in more CCSNe, bringing the number to a total of eight CCSNe with point-symmetric morphology.
I consider the point-symmetric morphologies of eight CCSNRs, supported by earlier studies of CCSNR morphologies, to be a step toward a breakthrough in determining the CCSN explosion mechanism. The JJEM cannot be ignored when analyzing CCSNe.
@@ 4.1 SN Ia scenarios
The diversity of normal and peculiar SNe Ia (e.g., DerKacy et al. 2023c; Desai et al. 2023b; Dwarkadas 2023b; Fausnaugh et al. 2023; Graur et al. 2023; Harris et al. 2023; Harvey et al. 2023; Hinse et al. 2023; Holland-Ashford et al. 2023; Huang et al. 2023; Keegans et al. 2023; Larison et al. 2023; Li et al. 2023b; Lu et al. 2023; Nascimento, França, & Reis 2023; O’Brien et al. 2023; Ogawa, Maeda, & Kawabata 2023; Pearson et al. 2023b; Peterson et al. 2023; Sharon & Kushnir 2023; Singh et al. 2023; Tiwari et al. 2023; Toy et al. 2023; Xi et al. 2023; Yang et al. 2023a; Bi, Woods, & Fabbro 2024; Giuffrida et al. 2024; Graur 2024; Pakmor et al. 2024; Petrecca et al. 2024; Singh et al. 2024; Weng et al. 2024; Williams et al. 2024) have motivated the consideration of all major SN Ia scenarios in 2023. I classify six scenarios as in Soker (2024a), and present them in Table 1 (for a different partition of the scenarios, in particular sub-classification of the SD scenario, see the table in the review by Liu, Röpke, & Han 2023).
@@
SN Ia Scenario
Core Degenerate
Double Degenerate - MED
Double Degenerate
Double Detonation
Single Degenerate
WD-WD collision
Name
CD
DD-MED
DD
DDet
SD-MED or SD
WWC
MED time
MED
MED
0
0
MED or 0
0
$\mathbf{[N_{\rm sur},M,Ej]}$${}^{[{\rm 2}]}$
$[0,M_{\rm Ch},{\rm S}]$
$[0,M_{\rm Ch},{\rm S}]$
[…]
$[1,M_{\rm Ch},{\rm S~{}or~{}N}]$
[…]
Point symmetry in the SNR
[…]
Very rare: A SNIP with point-symmetric PN.
Extremely rare.
Possible: a symbiotic progenitor; Low-mass CSM.
Extremely rare; large-scale departures from ellipticity.
Notes: (1) Cells in bold red indicate new suggestions/speculations from 2023 that I support in this review. The two SNRs Ia that I focus on are the peculiar SN 1181, which I suggest might result from the DD-MED or the CD scenarios, and SNR G1.9+0.3, which results from the CD scenario in a planetary nebula (SNIP). (2) Abbreviation. MED time: Merger to explosion delay time (includes mass transfer to explosion delay time). $N_{\rm exp}$ is the number of stars in the system at the time of the explosion. $N_{\rm sur}=1$ if a companion survives the explosion while $N_{\rm sur}=0$ if no companion survives the explosion; in some peculiar SNe Ia, the exploding WD is not destroyed and it also leaves a remnant, i.e., $N_{\rm sur}=2$. $M_{\rm Ch}$ for near-Chandrasekhar-mass explosion; sub-$M_{\rm Ch}$ for sub-Chandrasekhar mass explosion. The ejecta morphology Ej: S for scenarios that can lead to a spherical SNR while N indicates that the scenario cannot form a spherical SNR.
[…]
@@
In the double degenerate (DD) scenario (e.g., Dimitriadis et al. 2023; Karthik Yadavalli et al. 2023; Moran-Fraile et al. 2023; Qi, Liu, & Wang 2023; Srivastav et al. 2023a) and in the DD-MED scenario, two WDs merge after losing angular momentum to gravitational waves. In the DD scenario, the explosion occurs within a few dynamical timescales, like in the violent merger channel, before the binary system’s non-spherical imprints disappear. This DD scenario channel has no MED time, and the explosion is non-spherical. Practically there are two stars at the explosion, i.e., $N_{\rm exp}=2$. In the DD-MED scenario channel, the explosion occurs only after the merger remnant relaxes. It will likely be a spherical near-$M_{\rm Ch}$ explosion. At explosion, there is only one star (as wide companions play no roles), i.e., $N_{\rm exp}=1$.
@@
The double detonation (DDet) DDet scenario, where a helium layer accreted from a companion is ignited and detonates the WD, was relatively popular in 2023 (e.g., Collins et al. 2023; Zingale et al. 2023; Boos, Townsley, & Shen 2024) with attribution of this scenario to some specific peculiar SNe Ia (e.g., Liu et al. 2023b, c; Padilla Gonzalez et al. 2023a, b for four peculiar SNe Ia, respectively). Several studies attribute hyper-runaway/hypervelocity WDs to the WD helium-donor companion in the DDet scenario (e.g. El-Badry et al. 2023; Werner et al. 2023). However, new studies question whether this scenario can explain normal SNe Ia. Igoshev, Perets, & Hallakoun (2023) argue that the numbers and properties of hypervelocity WDs with velocities of $>1000{~{}\rm km}{~{}\rm s}^{-1}$ might at most account for a small fraction of SNe Ia, but might account for some peculiar SNe Ia. Braudo & Soker (2023) study the dynamics of the surviving WD and of the ejecta, and conclude that very fast runaway WDs require massive exploding CO WD. This makes such a scenario rare and supports the claim that the DDet scenario leads mostly to peculiar SNe Ia but not to normal SNe Ia.
@@
In the single degenerate (SD) scenario (e.g., Hosseinzadeh et al. 2023b; Li, Liu, & Wang 2023; Palicio et al. 2024, and Ruiz-Lapuente 2023 and Ruiz-Lapuente et al. 2023 for Tycho SNR) a WD accretes mass from a non-degenerate companion and explodes when reaching a near-Chandrasekhar mass. It can explode as it reaches the exploding mass, or there might be a long delay time from mass accretion to explosion (MED time). The non-detection of surviving companions in SNRs Ia, e.g., Shields et al. (2023) for SNR 0509-67.5, Tucker & Shappee (2023) for SN 2011fe, and Rodriguez, Yanza, & Dzib (2023) for Tycho’s SNR, continues to be a major challenge to the SD scenario (and also to the DDet scenario). Considering the difficulties of the SD scenario, I find it unjustified that in 2023 many papers mention only the SD and the DD scenarios, ignoring other SN Ia scenarios.
@@
Studies in 2023 have continued to explore both sub-Chandrasekhar (sub-$M_{\rm Ch}$) and near-Chandrasekhar ($M_{\rm Ch}$) explosions (e.g., Chakraborty et al. 2023; DerKacy et al. 2023a, b; Guo et al. 2023; Hoeflich et al. 2023; Leung & Nomoto 2023; Maeda et al. 2023b; Ni et al. 2023a; Yarbrough et al. 2023; Bravo, Isern, & Piersanti 2024).

### v4

@@ 3.2 The breakthrough: evidence for exploding jets
From Section 3.1, it is evident that although the delayed neutrino explosion mechanism is more popular in the literature than the JJEM, it nonetheless encounters some challenges. Here, I will argue that the point-symmetric morphologies of three CCSNRs that were studied in 2023 strongly support the JJEM and pose severe problems to the neutrino-driven explosion mechanism (for more point-symmetric CCSNe that were studied later in 2024, see Soker 2024e).
@@
Late observations, from months post-explosion before the nebular phase (e.g., Haynie & Piro 2023; Medler et al. 2023), through the nebular phase (e.g., DerKacy et al. 2023; Dessart et al. 2023a, b; Ertini et al. 2023; Fang & Maeda 2023; Liljegren et al. 2023; Margutti et al. 2023; Omand & Jerkstrand 2023; Ravi et al. 2023; Rizzo Smith, Kochanek, & Neustadt 2023; Wesson et al. 2023; Xiang et al. 2023, and then the remnant phase (e.g., Sato et al. 2023; Zsíros et al. 2024 and more below) can teach us about the progenitor properties and its environment at the explosion time and on the explosion mechanism. Before I describe evidence for jet-driven CCSNe by analyzing the CCSNR phase, I refer to the prediction of the delayed neutrino explosion mechanism that some massive stars collapse into a black hole with no explosion (or with only a very faint explosion). This collapse is termed ‘failed supernovae’.
@@
Some papers in 2023 continue to study this concept of a ‘failed supernova’ (e.g., Coughlin 2023; da Silva Schneider & O’Connor 2023; Disberg & Nelemans 2023; Kuroda & Shibata 2023; Li et al. 2023a; Paradiso et al. 2024). A ‘failed supernova’ is a prediction of the delayed neutrino explosion mechanism where neutrino heating cannot explode the star which in turn collapses to a black hole. On the other hand, in the JJEM, there are no failed CCSNe. Even if a black hole is formed, there is an explosion, and more likely, an energetic explosion (see review Soker 2022c). In the JJEM, a black hole is likely to be formed when the pre-collapse core is rapidly rotating because this leads to fixed-axis jets that are less efficient in exploding the core (Soker 2023b), i.e., an inefficient jet feedback mechanism (for a recent paper on the accretion disk around the central black hole in CCSNe see Fujibayashi et al. 2024). Soker (2023b) studies the continuum of explosion outcomes from zero pre-collapse core rotation to rapid pre-collapse rotation in the frame of the JJEM.
@@
The new study by Strotjohann, Ofek, & Gal-Yam (2023) and early results of Byrne & Fraser (2022) question the existence of a large population of ‘failed supernovae’. The detection of a black hole at the center of an SNR, as tentatively claimed by Balakrishnan et al. (2023) for SNR G351.9-0.9, shows that black holes can form in CCSNe, rather than in ‘failed supernovae’. The detection by Beasor et al. (2023) of a luminous infrared source at the position of the ‘failed supernova’ candidate N6946-BH1 shows, as they claim, that it better fits a type II intermediate luminosity optical transient (type II ILOT) than a ‘failed supernova’ (but see Kochanek 2024; Kochanek, Neustadt, & Stanek 2024). The above-cited studies challenge the prediction of the delayed neutrino explosion mechanism of ‘failed supernovae’. I reiterate my claim that studies of CCSNe that form black holes cannot ignore jets.
@@
Let me then turn to a prediction of the JJEM that in some (but not all) cases, the jittering jets imprint a point-symmetric morphology. Most of the jittering jets during the explosion process, typically lasting for about one to several seconds and more, deposit their entire energy in the core of the star and explode it. From that point on, the explosion is similar (but not identical) to that of the delayed-neutrino explosion mechanism. However, in some cases, the few last pairs of jets that are launched after the inner parts of the stars are exploded might leave an imprint on the outer ejecta. In most cases, only the last pair leaves an imprint, forming two opposite ‘ears’ (e.g., Soker 2023d). An ear is a protrusion from the main ejecta with a decreasing cross-section with distance from the center. In a minority of cases, more than one pair of jets might leave imprints, forming a point-symmetric morphology.
@@
In 2023, I identified point-symmetric morphologies in three CCSNRs. In Section 1 I discussed the point-symmetric morphology and why jets are most likely to shape such morphologies.
@@
In Soker (2024d), I studied the morphology of the CCSNR N63A based on new X-ray images that Karagöz et al. (2023) present. I present the image and my identification of a point-symmetric morphology in Figure 1. I identified the point-symmetric morphology to be the three pairs of ears that are marked in the figure. This point symmetry cannot be shaped by ejecta-CSM interaction or ejecta-ISM interaction (Section 1). Therefore, this point-symmetry strongly supports the JJEM. Moreover, the differences in sizes of the two ears in each pair are accounted for by the JJEM.
@@
In each pair, one ear is significantly larger than the opposite ear. This is not due to an asymmetrical medium on sides A and B (marked in the figure), as Ear 1a is larger than Ear 1b, but Ears 2a and 3a are smaller than Ears 2b and 3b, respectively. I, therefore, concluded (Soker, 2024d) that the asymmetry of each pair results from asymmetrical opposite jets at the time of the jets’ launching. This asymmetry likely results from the accretion disk that launches the jets in the JJEM having no time to relax fully. In the JJEM there are $\approx{\rm few}$ to $\approx 30$ jet-launching episodes, each lasting for $\approx 0.01-0.1{~{}\rm s}$. The viscous timescale of the accretion disk in each jet-launching episode is not much shorter than this timescale. Therefore, the disk has no time to fully relax to a thin accretion disk. Since the instabilities that must operate in the JJEM are likely to form each accretion disk with unequal sides of the temporarily equatorial plane, this asymmetry lasts for most of the jet-launching episode. The two unequal sides of the accretion disk are very likely to launch two opposite jets that have different energy, mass, and velocity. These asymmetrical two opposite jets in a pair will inflate unequal ears.
@@
In Soker (2023d), I classified 14 CCSNRs into five groups according to their morphology shaped by jets. I argued that the degree of pre-collapse core rotation is one of the main parameters determining the morphological class. The faster the pre-collapse core rotation is the less random the directions of the jets in the different jet-launching episodes are. For a very slow (or zero) pre-collapse core rotation, the jets jitter almost randomly. This, I suggested, might lead to a point-symmetric morphology with pairs of opposite structures in many directions. In that study, I identified a point-symmetric morphology in the Vela SNR. The Vela SNR is well-studied (e.g., Mayer et al. 2023; Ritchey 2023; Kargaltsev, Hare, & Lange 2024). I present the identification of the point-symmetric morphology in Figure 2 taken from Soker (2023d). Earlier studies already noted the existence of pairs of opposite structural features, as marked in the figure. However, only in Soker (2023d) did I explore the full point symmetry and connect it to jittering jets that exploded the progenitor of the Vela SNR (for newer images of Vela, see Mayer et al. 2023, and for newer analysis of its point symmetry, see Soker 2024e).
@@
Figure 2: The point-symmetric morphology of the Vela SNR, as identified by Soker (2023d). The ROSAT X-ray image is from Aschenbach et al. (1995), and the figure is adapted from figure 1 of Sapienza et al. (2021). Clumps A to F were identified by Aschenbach et al. 1995. The white AG-line and the labeling of the clumps are by Sapienza et al. (2021). The thick-yellow FJ-line and DE-line and the two dashed-black lines are marked by Soker (2023d). Each of the two dashed-black lines connects a clump to an assumed counter jet (for newer analysts, see Soker 2024e).
@@
SN 1987A has been one of the prime focuses of the SN community since its explosion (for 2023 papers see, e.g., Arendt et al. 2023; Cikota et al. 2023; Dohi et al. 2023; Fiorillo et al. 2023; Jones et al. 2023; Kangas et al. 2023; Larsson et al. 2023; Weishi Li et al. 2023; Bouchet et al. 2024; Chen, Chen, & Ono 2024b; Fransson et al. 2024; Ravi et al. 2024; Sapienza et al. 2024). I am interested in the ejecta morphology and its implications on the explosion mechanism. Ono et al. (2024) consider a bipolar explosion morphology for SN 1987A. They do not refer to jets, though.
@@
Refer to captionRefer to caption
Figure 3: An image of SN 1987A from JWST [source: NASA Webb Telescope Team, AUG 31, 2023; NASA, ESA, CSA: Matsuura, M. (Cardiff University), Arendt, R. (NASA’s Goddard Spaceflight Center & University of Maryland, Baltimore County), Fransson, C.]. [^3] For the original image see: https://www.nasa.gov/missions/webb/webb-reveals-new-structures-within-iconic-supernova/ Red represents emission at 4.44 microns, orange at 4.05 microns, yellow at 3.23 microns, cyan at 1.64 and 2.0 microns, and blue at 1.5 microns. The arrows in the left panel are additions from Soker (2024b). These arrows point at clumps, namely, some small areas in the outer regions of the blue-cyan structure that are brighter than their surroundings. The three black lines in the right panel, also from Soker (2024b), connect pairs of clumps that the arrows in the left panel point at. The red double-headed arrow indicates the direction of the elongated axis of the blue-cyan structure. The two red-dashed lines signify a possible structure of two opposite jets. (for newer images of SN 1987A, see Rosu et al. 2024, and for newer analysis of its point symmetry, see Soker 2024e).
Let me end with the study of the northeast jet of the SNR Cassiopeia A by Koo et al. (2023b). They conclude that the neutrino-driven explosion mechanism and the magnetocentrifugal jet-induced explosion (fixed-axis jets) have problems accounting for the properties of the northeast jet of Cassiopeia A. This is a known result. As I commented in the past, the JJEM can best explain the properties of Cassiopeia A (e.g., Soker 2017). New results (after this paper’s first and second versions were posted on the arXiv) by Bear & Soker (2024) show that Cassiopeia A also has a very robust point symmetry, as expected by the JJEM. An even later paper (Soker, 2024e) identifies point symmetry in more CCSNe, bringing the number to a total of eight CCSNe with point-symmetric morphology. That paper presents the images of these eight CCSNRs and compares them to the morphologies of the X-ray emitting gas in some groups and clusters of galaxies that are known to be shaped by jets. This comparison strongly suggests that jets shaped these CCSNRs as well. These point-symmetric CCSNRs are the CCSNR 0540-69.3, which has a point symmetry in its velocity map in a plane along the line of sight and a jet axis on the plane of the sky, and CCSNRs with point-symmetric morphologies on the plane of the sky, Cassiopeia A, N63A, Vela, SN 1987A, G321.3–3.9 which show clear point-symmetry, and CTB1 and G107.7-5.1 which show less prominent point-symmetry (tentative identification).
I consider the point-symmetric morphologies of eight CCSNRs, supported by earlier studies of CCSNR morphologies, to be a step toward a breakthrough in determining the CCSN explosion mechanism. A caveat is that there is not yet a quantitative classification of the point-symmetric morphology of CCSNRs. The JJEM cannot be ignored when analyzing CCSNe.
@@ 4.1 SN Ia scenarios
The diversity of normal and peculiar SNe Ia (e.g., DerKacy et al. 2023c; Desai et al. 2023b; Dwarkadas 2023b; Fausnaugh et al. 2023; Graur et al. 2023; Harris et al. 2023; Harvey et al. 2023; Hinse et al. 2023; Holland-Ashford et al. 2023; Huang et al. 2023; Keegans et al. 2023; Larison et al. 2024; Li et al. 2023b; Lu et al. 2023; Nascimento, França, & Reis 2023; O’Brien et al. 2023; Ogawa, Maeda, & Kawabata 2023; Pearson et al. 2024; Peterson et al. 2023; Sharon & Kushnir 2023; Singh et al. 2023; Tiwari et al. 2023; Toy et al. 2023; Xi et al. 2024; Yang et al. 2023a; Bi, Woods, & Fabbro 2024; Giuffrida et al. 2024; Graur 2024; Pakmor et al. 2024; Petrecca et al. 2024; Singh et al. 2024; Weng et al. 2024; Williams et al. 2024) have motivated the consideration of all major SN Ia scenarios in 2023. I classify six scenarios as in Soker (2024a) and present them in Table 1 (for a different partition of the scenarios, in particular sub-classification of the SD scenario, see the table in the review by Liu, Röpke, & Han 2023).
@@
Group $N_{\rm exp}=1$: Lonely WD $N_{\rm exp}=2$
Outcome Mostly normal SNe Ia Mostly peculiar SNe Ia
SN Ia Scenario Core Degenerate Double Degenerate - MED Double Degenerate Double Detonation Single Degenerate WD-WD collision
Name CD DD-MED DD DDet SD-MED or SD WWC
MED time MED MED 0 0 MED or 0 0
@@
Point symmetry in the SNR Expected in SNIPs with point-symmetric PNe. SN 1181; G1.9+0.3 Very rare: A SNIP with point-symmetric PN. SN 1181 Very rare: A SNIP with point-symmetric PN. Extremely rare. Possible: a symbiotic progenitor; Low-mass CSM. Extremely rare; large-scale departures from ellipticity.
Notes: (1) Cells in bold red indicate new suggestions/speculations from 2023 that I support in this review. The two SNRs Ia that I focus on are the peculiar SN 1181, which I suggest might result from the DD-MED or the CD scenarios, and SNR G1.9+0.3, which results from the CD scenario in a planetary nebula (SNIP). (2) Abbreviation. MED time: Merger to explosion delay time (includes mass transfer to explosion delay time). $N_{\rm exp}$ is the number of stars in the system at the time of the explosion. $N_{\rm sur}=1$ if a companion survives the explosion while $N_{\rm sur}=0$ if no companion survives the explosion; in some peculiar SNe Ia, the exploding WD is not destroyed, and it also leaves a remnant, i.e., $N_{\rm sur}=2$. $M_{\rm Ch}$ for near-Chandrasekhar-mass explosion; sub-$M_{\rm Ch}$ for sub-Chandrasekhar mass explosion. The ejecta morphology Ej: S indicates scenarios that can lead to a spherical SNR, while N indicates scenarios that cannot form a spherical SNR.
@@
In the double degenerate (DD) scenario (e.g., Dimitriadis et al. 2023; Karthik Yadavalli et al. 2023; Moran-Fraile et al. 2024; Qi, Liu, & Wang 2023; Srivastav et al. 2023a) and in the DD-MED scenario, two WDs merge after losing angular momentum to gravitational waves. In the DD scenario, the explosion occurs within a few dynamical timescales, like in the violent merger channel, before the binary system’s non-spherical imprints disappear. This DD scenario channel has no MED time, and the explosion is non-spherical. Practically, there are two stars at the explosion, i.e., $N_{\rm exp}=2$. In the DD-MED scenario channel, the explosion occurs only after the merger remnant relaxes. It will likely be a spherical near-$M_{\rm Ch}$ explosion. At explosion, there is only one star (as wide companions play no roles), i.e., $N_{\rm exp}=1$.
@@
The double detonation (DDet) DDet scenario, where a helium layer accreted from a companion is ignited and detonates the WD, was relatively popular in 2023 (e.g., Collins et al. 2023; Zingale et al. 2023; Boos, Townsley, & Shen 2024) with attribution of this scenario to some specific peculiar SNe Ia (e.g., Liu et al. 2023b, c; Padilla Gonzalez et al. 2023a, b for four peculiar SNe Ia, respectively). Several studies attribute hyper-runaway/hypervelocity WDs to the WD helium-donor companion in the DDet scenario (e.g. El-Badry et al. 2023; Werner et al. 2024). However, new studies question whether this scenario can explain normal SNe Ia. Igoshev, Perets, & Hallakoun (2023) argue that the numbers and properties of hypervelocity WDs with velocities of $>1000{~{}\rm km}{~{}\rm s}^{-1}$ might at most account for a small fraction of SNe Ia, but might account for some peculiar SNe Ia. Braudo & Soker (2024) study the dynamics of the surviving WD and the ejecta, and conclude that very fast runaway WDs require massive exploding CO WD. This makes such a scenario rare and supports the claim that the DDet scenario leads mostly to peculiar SNe Ia but not to normal SNe Ia.
@@
In the single degenerate (SD) scenario (e.g., Hosseinzadeh et al. 2023b; Li, Liu, & Wang 2023; Palicio et al. 2024, and Ruiz-Lapuente 2023 and Ruiz-Lapuente et al. 2023 for Tycho SNR) a WD accretes mass from a non-degenerate companion and explodes when reaching a near-Chandrasekhar mass. It can explode as it reaches the exploding mass, or there might be a long delay from mass accretion to the explosion (MED time). The non-detection of surviving companions in SNRs Ia, e.g., Shields et al. (2023) for SNR 0509-67.5, Tucker & Shappee (2024) for SN 2011fe, and Rodriguez, Yanza, & Dzib (2023) for Tycho’s SNR, continues to be a major challenge to the SD scenario (and also to the DDet scenario). Considering the difficulties of the SD scenario, I find it unjustified that in 2023 many papers mention only the SD and the DD scenarios, ignoring other SN Ia scenarios.
@@
Studies in 2023 have continued to explore both sub-Chandrasekhar (sub-$M_{\rm Ch}$) and near-Chandrasekhar ($M_{\rm Ch}$) explosions (e.g., Chakraborty et al. 2023; DerKacy et al. 2024, 2023; Guo et al. 2023; Hoeflich et al. 2023; Leung & Nomoto 2023; Maeda et al. 2023b; Ni et al. 2023a; Yarbrough et al. 2023; Bravo, Isern, & Piersanti 2024).